Solve by setting up the proper equation to describe the facts given and then carrying out the mathematical calculations to solve for the unknown variable(s).
The length of a rectangle is 3 times the width. The difference between the length and the width is 8 feet. Determine the rectangle’s width.

Answer by checkley77(12844)

L=3W
L-W=8
3W-W=8
2W=8
W=8/2
W=4 ANSWER FOR THE WIDTH.
L=3*4
L=12 ANSWER FOR THEW LENGTH.
PROOF:
12-4=8
8=8